# Towards Open Vocabulary Learning: A Survey

### v1

@@ 3.4 Open Vocabulary Segmentation
TABLE II: Representative works summarization and comparison in Sec. 3.5, Sec. 3.6 and Sec. 3.7. OV-VC refers to open vocabulary video classification. OV-3D refers to open vocabulary 3D recognition.
@@
Point Cloud Label (Base)
@@
3D Encoder (CLIP-vision distilled)
@@ 3.5 Open Vocabulary Video Understanding
In this section, we review open vocabulary learning on video tasks. Most works focus on designing the temporal fusion or association of VLM models in various settings, including action recognition and tracking.
@@
Video Classification. Traditional video classification methods usually require large datasets specific for video (e.g., Kinetics [231]). However, annotating video datasets requires very high costs. Using semantic information of label texts from web data may alleviate this deficiency. Building on the success of CLIP in image open vocabulary recognition, ActionCLIP [219] uses knowledge from image-based vision-language pre-training. It adds a temporal fusion layer for the zero-shot capability in video action recognition. A concurrent work I-VL [220] follows a similar paradigm by training the transformer layer on top of a frozen CLIP image encoder and also has a strong zero-shot capability. Another concurrent work, EVL [232], also employs the frozen CLIP for efficient video action recognition. X-CLIP [233] uses a video encoder that leverages the temporal information in the encoder for video recognition. It aligns the feature from the video encoder with the text encoder that is pre-trained on the image-based vision-language pairs (e.g., CLIP). To take advantage of other modalities, MOV [234] further fuses the audio information along with the pre-trained CLIP model to build a multi-modal open vocabulary video classification model. Text4Vis [235] focuses on the classifier and proposes to use the CLIP text encoder as the semantic target for video classification. Recently, Open-VCLIP [236] formulates the CLIP-to-video knowledge transfer as a continual learning problem and proposes Interpolated Weight Optimization to address the problem. AIM [237] tackles the open vocabulary video classification problem by adding adapt layers on top of the CLIP image encoder. ViFi-CLIP [221] reveals that a simple fine-tuning baseline (image-level feature extraction with CLIP visual encoder following temporal pooling) instead of advanced fusion layers can have a strong performance and further investigates the influence of prompt tuning. ASU [238] explores the use of fine-grained language features extracted by semantic units (e.g., head, arms, balloon, knee bend posture of exercise in the video) to guide the training of video classification. Recent advancement in open vocabulary video classification turns the attention to the language part of the VLM modeling. VicTR [239] introduces video-conditioned text representations to optimize the visual and text information jointly. MAXI [240] leverages LLMs to build a text bag for video without annotation by text expansion. The verbs, which are lacking in the realm of image, have the opportunity to participate in video-language modeling.
@@
Object Tracking and Video Instance Segmentation. The object tracking and video instance segmentation can also enjoy the rich knowledge of VLMs to build an open vocabulary tracker based on a close-set tracker. Going beyond the large vocabulary object tracking [241], to tackle the real-world multiple object tracking (MOT), OVTrack [222] first introduces large VLMs to the object tracking and tackles their proposed open vocabulary MOT (OV-MOT) task. Specifically, OVTrack extracts RoIs via an RPN and uses CLIP for knowledge distillation. A separate tracking head is used for tracking and supervised by pseudo LVIS videos. As for video instance segmentation (VIS), to make the VIS model capable of generalizing to novel classes in the real world, MindVLT [223] adopts a frozen CLIP backbone and proposes an end-to-end method with an open vocabulary classifier for segmenting and tracking unseen categories. It collects a large-vocabulary VIS dataset and tests their MindVLT on it. A concurrent work, OpenVIS [224], also tackles the open vocabulary video instance segmentation task but in a different manner. It generates the class-agnostic mask of instances and leverages the masks to crop raw images to feed into the CLIP visual encoder for calculating class scores.
3.6 Open Vocabulary 3D Scene Understanding
@@
In this section, we review the open vocabulary approaches that are used in 3D scene understanding tasks. We mainly survey the works for point cloud classification and segmentation, which explores the knowledge of 2D VLMs into 3D.
@@
3D to 2D projection. Extending the Open Vocabulary mechanism into 3D perception is not trivial since it is hard to collect enough data for point-language contrastive training like what CLIP does in 2D perception. To mitigate the gap, PointCLIP [225] takes the first step to 3D open vocabulary perception by projecting the 3D point cloud to a 2D plane and extracting visual features from the projected depth map via CLIP visual encoder. In this way, the knowledge from the pre-trained CLIP can be put to good use. Rather than directly using the CLIP visual encoder for visual feature extraction of depth map, CLIP2Point [242] aligns the RGB image feature from the CLIP visual encoder and depth feature from a depth encoder through contrastive learning. Then, the depth feature can be aligned with the language embedding. PointCLIPV2 [243] further advances depth map generation and 3D prompting to get a significant improvement.
3D and 2D Alignment. Projecting 3D points to 2D may not take full use of 3D information. Accordingly, beyond the 3D to 2D projection methods, ULIP [226] proposes to align the 3D feature extracted by the 3D backbone to the CLIP-aligned visual and text feature to enable the zero-shot capability. In the driving scene, LidarCLIP [244] also aligns the 3D and 2D features, which is applicable for 3D open vocabulary scene understanding and lidar-to-image generation.
Downstream Tasks. 3D object detection and segmentation also demand to identify unknown objects in 3D scene. Inspired by the success of 2D open vocabulary detection [160], aiming at taking full use of both image and point cloud modalities for 3D detection, OV-3DETIC [227] and OV-3DET [245] take a further step to decouple the localization and recognition into two modalities: the point cloud is taken into account for localization, while RGB image is used for recognition. Sharing similar motivation that there lacks 3D-text pairs for contrastive training, PLA [229] focuses on the 3D scene understanding. It uses multi-view images to extract visual features for 3D encoder contrastive learning to enable the open vocabulary capability. Meanwhile, OpenMask3D [246] presents a multi-view fusion approach for open vocabulary 3D instance segmentation. With the aid of 3D class-agnostic masks from point cloud inputs, the model aggregates per-mask CLIP features, which achieves good zero-shot ability in instance classification heads. More recent work, CLIP-FO3D [247], focuses on 3D semantic segmentation and uses visual feature distillation for 3D feature learning to directly transfer CLIP’s knowledge and avoid extra annotation. In the driving scene, to mitigate the heavy dependence on the point cloud data annotation and fast generalization to the new scenes, CLIP2Scene [248] proposes to train a 3D network via semantic-driven cross-modal contrastive learning for 3D segmentation.
@@ 3.7 Closely Related Tasks
In this section, we add several closely related tasks to augment open vocabulary learning. We survey the following aspects, including class agnostic detection and segmentation, open world object detection, and open-set panoptic segmentation.
@@
$\bullet$ Class Agnostic Detection and Segmentation. The goal of class agnostic detection and segmentation is to learn a general region proposal system that can be used in different scenes. For detection, OLN [249] replaces the binary classification head in RPN by predicting the IoU score of foreground objects, which proves the generalization ability on cross-dataset testing. For example, the model trained on COCO shows the localization ability on the LVIS dataset. Open world instance segmentation [250] aims to correctly detect and segment instances whether their categories are in training taxonomy or not. GGN [251] leverages the bottom-up grouping idea, which combines a local measure of pixel affinity with instance-level mask supervision, producing a training regimen designed to make the model more generalizable. Meanwhile, from the data augmentation view, MViT [252] combines the Deformable DETR and CLIP text encoder for the open world class-agnostic detection, where the authors build a large dataset by mixing existing detection datasets. For the segmentation domain, Entity segmentation [253, 254] aims to segment all visual entities without predicting their semantic labels. The goal of such segmentation is to obtain high-quality and generalized segmentation results. Fined-grained entity [254] presents a two-view image crop ensemble method named CropFormer to enhance the fine-grained details. Recently, SAM [33] proposes more generalized prompting methods, including masks, points, boxes, and texts. In particular, the authors build a larger dataset with 1 billion masks following the spirits of CLIP. The SAM achieves zero-shot testing in various segmentation datasets, which are extremely generalizable.
@@
$\bullet$ Open World Object Detection. Open world recognition [27] requires the model to identify novel categories and label them as ‘unknown.’ Then, the novel categories are progressively annotated. The model learns incrementally with the data and recognizes the newly-annotated categories. Inspired by this setting, open world object detection (OWOD) [255] expands the recognition task to the detection setting. To tackle the proposed problem, the authors propose a novel method that utilizes contrastive clustering and an energy-based unknown identification module. Following this work, OW-DETR [256] introduces an end-to-end transformer-based framework. It consists of three dedicated components, namely, attention-driven pseudo-labeling, novelty classification, and objectness scoring, to explicitly address the OWOD challenge. Meanwhile, open world DETR [257] proposes a two-stage training approach based on Deformable DETR. It focuses on alleviating catastrophic forgetting when the annotations of the unknown classes become available incrementally by using knowledge distillation and exemplar replay technologies. PROB [258] proposes a probabilistic objectness head into the OW-DETR to better mine unknown background classes. Recently, researchers [259] propose a new setting, named unknown-classified open world object detection (UC-OWOD), which aims to detect unknown instances and classify them into different unknown classes. This task is close to the open vocabulary object detection but requires incremental learning.
@@
$\bullet$ Open-Set Panoptic Segmentation. Similar to other open-set tasks, open-set panoptic segmentation (OSPS) [260] requires the model to identify novel categories as ‘unknown’ in a panoptic segmentation task. The ’unknown’ classes are chosen from the thing classes (foreground objects). The authors apply exemplar theory for novel class discovery. After that, Dual [261] proposes a divide-and-conquer scheme to develop a dual decision process for OSPS. The results indicate that by properly combining a known class discriminator with an additional class-agnostic object prediction head, the OSPS performance can be significantly improved.
4 Benchmark Results

### v4

@@ 3.4 Open Vocabulary Segmentation
TABLE IV: Representative works summarization and comparison in Sec. 3.5, Sec. 3.6 and Sec. 3.7. OV-VC refers to open vocabulary video classification. OV-3D refers to open vocabulary 3D recognition.
@@
PointCLIPV2 [274]
OV-3D
None
CLIP-text + GPT-3
CLIP-vision (projected depth map)
PointCLIPV2 generates prompts with LLMs to enhance performance.
@@
Point Cloud Label
@@
Point Cloud Encoder + CLIP-vision
@@
CLIP2 [275]
OV-3D
None
CLIP-text
Point Cloud Encoder + CLIP-vision
CLIP2 leverages real-world point cloud data and CLIP to train the 3D encoder for 3D open vocabulary segmentation.
OpenShape [276]
OV-3D
Point Cloud Label
CLIP-text + GPT-4
Point Cloud Encoder + CLIP-vision
OpenShape builds a large-scale dataset that provides image, point cloud, and text triplets.
@@
OpenScene [116]
OV-3D-SS
None
CLIP-text
3D Encoder + LSeg [14]
OpenScene train a 3D Encoder yielding dense features co-embedded with text and image pixels for open vocabulary semantic segmentation.
@@ 3.5 Open Vocabulary Video Understanding
We also review several video open vocabulary tasks. Most works focus on designing VLM models’ temporal fusion or association in various settings, including action recognition and tracking.
@@
Video Classification. Traditional video classification methods usually require large datasets specific to video (e.g., Kinetics [280]). However, annotating video datasets requires very high costs. Using semantic information of label texts from web data may alleviate this deficiency. Building on the success of CLIP in image open vocabulary recognition, ActionCLIP [268] uses knowledge from image-based vision-language pre-training. It adds a temporal fusion layer for the zero-shot capability in video action recognition. A concurrent work I-VL [269] follows a similar paradigm by training the transformer layer on top of a frozen CLIP image encoder and also has a strong zero-shot capability. Another concurrent work, EVL [281], also employs the frozen CLIP for efficient video action recognition. X-CLIP [282] uses a video encoder that leverages the temporal information in the encoder for video recognition. It aligns the feature from the video encoder with the text encoder that is pre-trained on the image-based vision-language pairs (e.g., CLIP). To take advantage of other modalities, MOV [283] further fuses the audio information and the pre-trained CLIP model to build a multi-modal open vocabulary video classification model. Recently, Open-VCLIP [284] formulates the CLIP-to-video knowledge transfer as a continual learning problem and proposes Interpolated Weight Optimization to address the issue. AIM [285] tackles the open vocabulary video classification problem by adding adapt layers on top of the CLIP image encoder. ViFi-CLIP [270] reveals that a simple fine-tuning baseline (image-level feature extraction with CLIP visual encoder following temporal pooling) instead of advanced fusion layers can have a strong performance and further investigates the influence of prompt tuning. ASU [286] explores the use of fine-grained language features extracted by semantic units (e.g., head, arms, balloon, knee bend posture of exercise in the video) to guide the training of video classification. Recent advancements in open vocabulary video classification have turned our attention to the language part of VLM modeling. VicTR [287] introduces video-conditioned text representations to optimize the visual and text information jointly. MAXI [123] leverages LLMs to build a text bag for video without annotation by text expansion. The verbs, which are lacking in the realm of image, have the opportunity to participate in video-language modeling.
@@
Object Tracking and Video Instance Segmentation. The object tracking and video instance segmentation can also enjoy the rich knowledge of VLMs to build an open vocabulary tracker based on a close-set tracker. Going beyond the large vocabulary object tracking [288], to tackle the real-world multiple object tracking (MOT), OVTrack [271] first introduces large VLMs to the object tracking and tackles their proposed open vocabulary MOT (OV-MOT) task. Specifically, OVTrack extracts RoIs via an RPN and uses CLIP for knowledge distillation. A separate tracking head is used for tracking and supervised by pseudo-LVIS videos. As for video instance segmentation (VIS), to make the VIS model capable of generalizing to novel classes in the real world, MindVLT [57] adopts a frozen CLIP backbone and proposes an end-to-end method with an open vocabulary classifier for segmenting and tracking unseen categories. It collects a large-vocabulary VIS dataset and tests their MindVLT on it. A concurrent work, OpenVIS [272], also tackles the open vocabulary video instance segmentation task but in a different manner. It generates the class-agnostic mask of instances and leverages the masks to crop raw images to feed into the CLIP visual encoder for calculating class scores. Beyond the open vocabulary video instance segmentation, DVIS++ [289] proposes the first open vocabulary universal video segmentation scheme supporting video semantic segmentation, video instance segmentation, and video panoptic segmentation. DVIS++ [289] leverages the knowledge from the CLIP backbone and adopts mask pooling on the CLIP backbone for enabling open vocabulary capability.
3.6 Open Vocabulary 3D Understanding
@@
In this section, we review the open vocabulary approaches used in 3D scene understanding tasks. We mainly survey the works for point cloud classification and segmentation, which explores the knowledge of 2D VLMs into 3D.
@@
3D Recognition. The VLM has shown great success in the zero-shot or few-shot learning of 2D images by leveraging a huge amount of image data. However, such internet-scale data is not available for the 3D point cloud. To this extent, extending the open vocabulary mechanism into 3D perception is not trivial since it is hard to collect enough data for point-language contrastive training like what CLIP does in 2D perception. To mitigate the gap, PointCLIP [273] takes the first step to 3D open vocabulary perception. The principle insight behind PointCLIP is that the 3D point cloud can be converted to CLIP-recognizable images. By projecting the 3D point cloud to a 2D plane and extracting visual features from the projected depth map via the CLIP visual encoder, the point cloud feature can be naturally aligned with the language feature extracted by the language encoder. Then, the whole framework is with 3D point cloud zero-shot capability, just like 2D images. However, simply projecting the point cloud to depth maps may yield inferior performance. To further improve the performance, rather than directly using the CLIP visual encoder for visual feature extraction of depth map, CLIP2Point [290] aligns the RGB image feature from the CLIP visual encoder and depth feature from a depth encoder through contrastive learning. Specifically, they collect image-depth pairs to align the from-scratch depth map encoder with the CLIP image encoder at training time and only use the depth map encoder at inference time. Then, the depth feature can be aligned with the language embedding, facilitating the 3D point cloud zero-shot capability with CLIP. Also aiming at improving the performance of PointCLIP [273], PointCLIPV2 [274] focuses on different perspectives and proposes a realistic shape projection scheme along with an LLM-based 3D prompting generation scheme. Although without extra annotations, PointCLIPV2 achieves a significant improvement over PointCLIP [273].
Though projecting the point cloud to depth maps can make it easy to leverage the existing 2D visual encoders pre-trained in the CLIP, it may not exhaustively use the 3D information in the point cloud. Accordingly, beyond the depth map-based methods, ULIP [117] first collects multi-modalities triplets (point cloud, image, and text) for training the 3D backbone. Since CLIP already aligns the language and image encoders, it only needs to align the 3D backbone to the image-language feature space. With only small-scale data, ULIP [117] aligns the 3D feature extracted by the 3D backbone to the CLIP-aligned visual and text feature to enable the zero-shot capability and enhance the standard 3D recognition capability. The advantage of ULIP [117] also includes unifying the three modalities, which may bring more downstream applications, such as image-to-3D retrieval. Though ULIP [117] aligns the 3D and 2D and thus enables the native 3D open vocabulary capability, it trains on a small-scale dataset, which limits its performance. CLIP2 [275] resorts to finding training samples from the real world and proposes Triplet Proxies Collection scheme for finding instance-level 3D point cloud, 2D image crop, and text triplets for 3D recognition. OpenShape [276] scales up both the dataset and backbone. For the dataset, they build a text-3D shape dataset containing 876k training shapes with over 1k categories. They also explore adopting larger 3D backbones and their performance. As a result, OpenShape [276] drastically enhances the performance on 3D zero-shot capability. LidarCLIP [291] also aligns the 3D and 2D features with a similar approach, but LidarCLIP [291] focuses on the driving scenes. Besides the 3D zero-shot capability, LidarCLIP [291] also shows point cloud captioning and lidar-to-image generation capability with off-the-shelf foundation models.
3D Object Detection. In 3D object detection, the training data is often hard to get, and existing 3D detectors are often trained on very limited classes, which means they are hard to generalize to novel classes. Inspired by the success of 2D open vocabulary detection [134], aiming at taking full use of both image and point cloud modalities for 3D detection, OV-3DETIC [277] and OV-3DET [292] proposes to decouple the localization and recognition in point cloud object detection into localization and recognition. For localization, 2D pre-trained detectors can be adapted to train the 3D detector by back-projecting the 2D bounding box and further optimizing it by the point cloud. For recognition, they propose aligning the regional features of 3D and 2D encoders with the corresponding text feature. Therefore, during the inference, 3DETIC [277] and OV-3DET [292] can generalize to unseen classes with the knowledge in the CLIP. While OV-3DET [292] achieves significant performance on the Open-Vocabulary 3D detection, it relies on the localization capability of pre-trained 2D detectors, which may limit the 3D object discovery capability. To solve this problem, CODA [293] proposes to discover objects with both 2D semantic prior and 3D geometry prior, and further proposes a cross-modal alignment module to align 3D and 2D features. Also trying to get rid of the 2D image detector, concurrent work Object2Scene [294] proposes to augment the existing 3D object detection datasets with large-scale 3D object datasets. It also introduces a new framework, L3Det [294], for 3D object-text alignment. As a result, both CODA [293] and L3Det [294] outperforms OV-3DET [292]. As driving scenarios may yield new challenges, the above-mentioned methods may have an inferior performance. OpenSight [295] focuses on the outdoor scenes and proposes to leverage the knowledge from 2D grounding DINO [214] to train the 3D detectors. In particular, it uses the size prior (e.g., cars are 1.96m wide) generated by the LLM to re-calibrate the output of the model to make the model more suitable for the driving scene.
3D Scene Understanding. In the understanding of the 3D scene, the same obstacle, lacking training data, also limits the capability of generalization. Sharing a similar motivation that there is a lack of 3D-text pairs for point-language contrastive training, PLA [124] proposes to extract features from multi-view images sampled from a scene to generate descriptions with pre-trained language models. The descriptions can then be used to extract language features to train 3D backbone-language alignment. With the language-aligned 3D backbone, it is natural to conduct open vocabulary segmentation like with the 2D image. OpenScene [116] proposes to link the 3D features for each point with CLIP features for each pixel. By back-projecting the 2D pixels to 3D space, each point in the point cloud can be ensembled with features of several pixels in different views. The 3D-text co-embedding makes it possible for open vocabulary semantic segmentation. Inspired by MaskCLIP [120], a concurrent work CLIP-FO3D [296] also focuses on 3D semantic segmentation and has a similar approach. It leverages the feature map extracted by the CLIP visual encoder and directly transfers CLIP’s knowledge without any extra annotation. A recent work, RegionPLC [297] on 3D open vocabulary semantic segmentation, considers the benefits of both PLA [124] and OpenScene [116]. It proposes to use the dense caption of any random region to extract more information in CLIP for 3D backbone distillation. Fine-grained dense supervision in RegionPLC [297] beyond view-level or instance-level further improves performance compared to PLA [124]. Different from training 3D backbones like in [124, 116, 296, 297], PartSLIP [298] and SATR [299] directly projecting the 3D point cloud and mesh to 2D plane and uses GLIP [231] to localize and segment. Meanwhile, OpenMask3D [300] mainly focuses on the 3D instance segmentation. It first generates class-agnostic instance mask proposals with the 3D point cloud. Then, OpenMask3D [300] selects views and projects 3D masks into 2D images. The 2D masks are further refined by SAM [42]. Finally, the 2D masks can be fed into the CLIP visual encoder to generate label prediction with the help of the CLIP language encoder. While OpenMask3D [300] has a strong performance, it requires point cloud-2D image pairs for training. On the contrary, OpenIns3D [301] does not require 2D images but only needs point clouds with colors. OpenIns3D [301] proposes a ”Mask-Snap-Lookup” pipeline that first learns class-agnostic mask proposals, then generates synthetic scene-level images, and finally assigns categories for each proposal. Another work trying to improve OpenMask3D [300] is Open3DIS [302]. Open3DIS [302] improves the 3D mask proposal quality by employing a 3D instance network and a 2D-guide-3D Instance Proposal Module. After the mask proposal, it aggregates CLIP features in a multi-scale, multi-view manner. Open3DIS [302] achieves a stronger performance over OpenMask3D [300] on several datasets. In the driving scene, to mitigate the heavy dependence on the point cloud data annotation and fast generalization to the new scenes, CLIP2Scene [303] proposes to train a 3D network via semantic-driven cross-modal contrastive learning for 3D segmentation. CLIP2Scene [303] also considers spatial-temporal semantic consistency regularization to facilitate training.
@@ 3.7 Closely Related Tasks
$\bullet$ Class Agnostic Detection and Segmentation. The goal of class agnostic detection and segmentation is to learn a general region proposal system that can be used in different scenes. For detection, OLN [304] replaces the binary classification head in RPN by predicting the IoU score of foreground objects, which proves the generalization ability on cross-dataset testing. For example, the model trained on COCO shows the localization ability on the LVIS dataset. Open world instance segmentation [305] aims to correctly detect and segment instances whether their categories are in training taxonomy. GGN [306] leverages the bottom-up grouping idea, combining a local pixel affinity measure with instance-level mask supervision, producing a training regimen designed to make the model more generalizable. Meanwhile, from the data augmentation view, MViT [307] combines the Deformable DETR and CLIP text encoder for the open world class-agnostic detection, where the authors build a large dataset by mixing existing detection datasets. For the segmentation domain, Entity segmentation [308, 309] aims to segment all visual entities without predicting their semantic labels. The goal of such segmentation is to obtain high-quality and generalized segmentation results. Fined-grained entity [309] presents a two-view image crop ensemble method named CropFormer to enhance the fine-grained details. Recently, SAM [42] proposes more generalized prompting methods, including masks, points, boxes, and texts. In particular, the authors build a larger dataset with 1 billion masks following the spirits of CLIP. The SAM achieves zero-shot testing in various segmentation datasets, which are highly generalizable.
@@
$\bullet$ Open World Object Detection. Open world recognition [27] requires the model to identify novel categories and label them as “unknown.” Then, the novel categories are progressively annotated. The model learns incrementally with new data and recognizes the newly-annotated categories. Inspired by this setting, open world object detection (OWOD) [32] expands the recognition task to object detection. The authors propose a novel method that utilizes contrastive clustering and an energy-based unknown identification module. Following them, OW-DETR [310] introduces an end-to-end transformer-based framework. It consists of three dedicated components, namely, attention-driven pseudo-labeling, novelty classification, and objectness scoring, to address the OWOD challenge explicitly. Meanwhile, open world DETR [311] proposes a two-stage training approach based on Deformable DETR. It focuses on alleviating catastrophic forgetting when the annotations of the unknown classes become available incrementally using knowledge distillation and exemplar replay technologies. PROB [312] proposes a probabilistic objectness head into the OW-DETR to better mine unknown background classes. Recently, researchers [313] propose unknown-classified open world object detection (UC-OWOD), which aims to detect and classify unknown instances into different unknown classes. This task is close to the open vocabulary object detection but requires incremental learning.
@@
$\bullet$ Open-Set Panoptic Segmentation. Similar to other open-set tasks, open-set panoptic segmentation (OSPS) [314] requires the model to identify novel categories as ‘unknown’ in a panoptic segmentation task. The ’un classes are chosen from the thing classes (foreground objects). The authors apply exemplar theory for novel class discovery. After that, Dual [315] proposes a divide-and-conquer scheme to develop a dual decision process for OSPS. The results indicate that by properly combining a known class discriminator with an additional class-agnostic object prediction head, the OSPS performance can be significantly improved.
4 CHALLENGES AND OUTLOOK
4.1 Challenges
Base Classes Over-fitting Issues. Most approaches detect and segment novel objects by learning proposals from base class annotations. Thus, there are natural gaps in shapes and semantic information between novel and base objects. VLM models can bridge such gaps via pre-trained visual-text knowledge. However, most detectors still easily overfit the base classes when the novel classes have similar shapes and semantics, since these classes are trained with higher confidence scores. More fine-grained feature discriminative modeling [316, 317, 318], including parts or attributes, is required to handle these problems.
Training Costs. Most state-of-the-art methods need huge data for pre-training to achieve good performances. However, the costs are expensive and even unavailable for many research groups to follow. Thus, with the aid of VLM, designing more efficient data learning pipelines or learning methods is more practical and affordable. One simple solution can be adopting a frozen backbone. However, this may limit the representation capacity.
Across Dataset Generation and Evaluation. As shown in the benchmark section, current state-of-the-art methods design specific models for each benchmark. Designing one shared model [214, 319, 259, 320, 321] across different datasets on open vocabulary detection and segmentation follows the origin spirits of open vocabulary learning. However, there are still performance gaps between unified models and dataset-specific models in several OVD benchmarks, such as LVIS datasets.
Better Benchmarks and Metrics. Since several classes contain overlapping concepts (for example, building and tower, person and woman), designing more new metrics [322] is also needed to better measure the open vocabulary methods. Moreover, current datasets are still small. To realize real open vocabulary settings, more datasets like SAM-1B [42] are needed.
4.2 Future Work
Explore Foundation Models With Specific Adapter For Custom Tasks. Vision foundation models [42, 15] can achieve good zero-shot performance on several standard classification and segmentation datasets. However, for several custom tasks, such as medical image analysis, and aerial images, there are still many corner cases. Thus, designing a task-specific adapter is needed for these custom tasks. Such adapters can fully utilize the knowledge of pre-trained foundation models. One possible solution is to explore the in-context learning [323, 324] to fully explore or connect the knowledge of VLMs and LLMs.
Combining with Large Language Models. Compared with VLMs, most LLMs contain more text concepts, which naturally have a broader scope than various dataset taxonomies, even larger than the recent V3Det [50] dataset. Thus, how to better align the LLMs knowledge [325] with visual detectors or segmenters to achieve stronger zero-shot results still needs exploration.
5 Conclusion
This survey offers a detailed examination of the latest developments in open vocabulary learning in computer vision, which appears to be a first of its kind. We provide an overview of the necessary background knowledge, which includes fundamental concepts and introductory knowledge of detection, segmentation, and vision language pre-training. Following that, we summarize more than 50 different models used for various scene understanding tasks. For each task, we categorize the methods based on their technical viewpoint. Additionally, we provide information regarding several closely related domains. In the experiment section, we provide a detailed description of the settings and compare results fairly. Finally, we summarize several challenges and also point out several future research directions for open vocabulary learning.
Acknowledgement. This work is supported by the National Key Research and Development Program of China (No. 2023YFC3807600) and the interdisciplinary doctoral grants (iDoc 2021-360) from the Personalized Health and Related Technologies (PHRT) of the ETH domain.
TABLE V: Open vocabulary video classification results. The mean and standard deviation of results are reported.
Method Vision Backbone Frames Ground Truth UCF101 HMDB51 K600 (Top-1) K600 (Top-5) ActionCLIP [268](Arxiv ’2109) ViT-B/16 32 58.3 $\pm$ 3.4 40.8 $\pm$ 5.4 66.7 $\pm$ 1.1 91.6 $\pm$ 0.3 I-VL [269](ECCV ’22) ViT-B/16 32 69.3 $\pm$ 4.2 44.3 $\pm$ 2.2 55.8 $\pm$ 0.7 81.4 $\pm$ 0.3 X-CLIP [282](ECCV ’22) ViT-B/16 32 72.0 $\pm$ 2.3 44.6 $\pm$ 5.2 65.2 $\pm$ 0.4 86.1 $\pm$ 0.8 Text4Vis [326](AAAI ’23) ViT-L/14 16 - - 68.9 $\pm$ 1.0 - ViFi-CLIP [270](CVPR ’23) ViT-B/16 32 76.8 $\pm$ 0.7 51.3 $\pm$ 0.6 71.2 $\pm$ 1.0 92.2 $\pm$ 0.3 ASU [286](Arxiv ’2303) ViT-B/16 32 75.0 $\pm$ 3.7 48.1 $\pm$ 2.8 67.6 $\pm$ 0.2 87.2 $\pm$ 0.3 MAXI [123](Arxiv ’2303) ViT-B/16 32 78.2 $\pm$ 0.8 52.3 $\pm$ 0.7 71.5 $\pm$ 0.8 92.5 $\pm$ 0.4 VicTR [287](Arxiv ’2304) ViT-B/16 32 72.4 $\pm$ 0.3 51.0 $\pm$ 1.3 - Open-VCLIP [284](ICML ’23) ViT-B/16 16 83.4 $\pm$ 1.2 53.9 $\pm$ 1.2 73.0 $\pm$ 0.8 93.2 $\pm$ 0.1 Open-VCLIP [284](ICML ’23) ViT-L/14 16 87.6 $\pm$ 1.2 59.0 $\pm$ 0.6 81.1 $\pm$ 0.8 96.3 $\pm$ 0.3
TABLE VI: The performance comparison on LV-VIS [57] validation and test set. The $AP_{all}$, $AP_{base}$, and $AP_{novel}$ mean the average precision of overall categories, base categories, and novel categories.
Method Backbone Val Test
$mAP_{all}$ $mAP_{base}$ $mAP_{novel}$ $mAP_{all}$ $mAP_{base}$ $mAP_{novel}$
DetPro [217]-SORT [327] R50 9.9 14.9 4.1 5.7 8.8 2.8
Detic [134]-SORT [327] R50 10.1 15.1 4.2 5.7 8.5 3.1
DetPro [217]-OWTB [328] R50 11.8 17.1 5.7 7.1 9.9 4.0
Detic [134]-OWTB [328] R50 11.5 16.8 5.3 7.1 10.1 3.8
MindVLT [57] R50 14.1 18.1 9.4 8.6 10.7 6.4
Detic [134]-SORT [327] SwinB 14.9 20.7 8.2 11.0 16.2 5.4
Detic [134]-OWTB [328] SwinB 15.8 21.5 9.1 11.7 16.4 6.6
MindVLT [57] SwinB 21.4 26.3 15.8 14.9 18.8 10.6
TABLE VII: The performance comparision of zero-shot classification on Objaverse-LVIS [329], ModelNet40 [330], and ScanObjectNN [331].
Method Training Objaverse-LVIS [329] ModelNet40 [330] ScanObjectNN [331]
Top1 Top3 Top5 Top1 Top3 Top5 Top1 Top3 Top5
PointCLIP [273] None 1.9 4.1 5.8 19.3 28.6 34.8 10.5 20.8 30.6
PointCLIP v2 [274] 4.7 9.5 12.9 63.6 77.9 85.0 42.2 63.3 74.5
CG3D [332] ShapeNet [333] 5.0 9.5 11.6 48.7 60.7 66.5 42.5 57.3 60.8
CLIP2Point [290] 2.7 5.8 7.9 49.5 71.3 81.2 25.5 44.6 59.4
ULIP-PointBERT [117] 6.2 13.6 17.9 60.4 79.0 84.4 51.5 71.1 80.2
OpenShape-SparseConv 11.6 21.8 27.1 72.9 87.2 93.0 52.7 72.7 83.6
OpenShape-PointBERT 10.8 20.2 25.0 70.3 86.9 91.3 51.3 69.4 78.4
ULIP-PointBERT [276] OpenShape [276] 21.4 38.1 46.0 71.4 84.4 89.2 46.0 66.1 76.4
OpenShape-SparseConv [276] (w/o Objaverse-LVIS [329]) 37.0 58.4 66.9 82.6 95.0 97.5 54.9 76.8 87.0
OpenShape-PointBERT [276] 39.1 60.8 68.9 85.3 96.2 97.4 47.2 72.4 84.7
ULIP-PointBERT [276] OpenShape [276] 26.8 44.8 52.6 75.1 88.1 93.2 51.6 72.5 82.3
OpenShape-SparseConv [276] 43.4 64.8 72.4 83.4 95.6 97.8 56.7 78.9 88.6
OpenShape-PointBERT [276] 46.8 69.1 77.0 84.4 96.5 98.0 52.2 79.7 88.7
[…]
Method ScanNet [334] nuScenes [335]
B15/N4 B12/N7 B10/N9 B12/N3 B10/N5
LSeg-3D [124] 00.0 / 64.4 / 00.0 00.9 / 55.7 / 00.1 01.8 / 68.4 / 00.9 00.6 / 74.4 / 00.3 0.00 / 71.5 / 0.00
PLA w/o training [124] 39.7 / 68.3 / 28.0 24.5 / 70.0 / 14.8 25.7 / 75.6 / 15.5 25.5 / 75.8 / 15.4 10.7 / 76.0 / 05.7
PLA [124] 65.3 / 68.3 / 62.4 55.3 / 69.5 / 45.9 53.1 / 76.2 / 40.8 47.7 / 73.4 / 35.4 24.3 / 73.1 / 14.5
RegionPLC [297] 69.9 / 68.4 / 71.5 65.1 / 69.6 / 61.1 58.8 / 76.6 / 47.7 62.0 / 75.8 / 52.4 36.6 / 76.7 / 24.1
Fully-Sup (Upper Bound) [297] 73.3 / 68.4 / 79.1 70.6 / 70.0 / 71.8 69.9 / 75.8 / 64.9 73.7 / 76.6 / 71.1 74.8 / 76.8 / 72.8
Appendix A Benchmark Results